# Minimal cellular automaton model with heterogeneous cell sizes predicts epithelial colony growth

### v1

@@ II Parameter calibration
Figure 3: Asymmetry $G(f)$ of the daughter cell sizes after cell division is well described by a Gaussian distribution with expectation value $\bar{f}=1/2$, standard deviation $\sigma=1/8$, and a cut-off at large and small cell sizes $p\in[1/4,3/4]$ (black dots represent data from Ref. (PulHufNevStrSigFygShr2012, , Fig. 4C (inset)) obtained from tracking single-cell divisions, Gaussian distribution shown as red line). It is assumed that during division of a cell, one daughter cell has a fraction $f$ of the size of the mother cell, while the other cell has a fraction $(1-f)$. The probability distribution $G(f)$ of these fractions is normalized to the domain $f\in[1/4,3/4]$, reflecting the absence of smaller or larger daughter cells.
@@
The biologically motivated parameters $\gamma(A)$, $v_{\text{max}}$, $\alpha$ of the CA model are calibrated based on in-vitro observations, mostly of single-cell dynamics, during the growth of normally differentiating epithelial cell cultures of Madin-Darby canine kidney cells PulHufNevStrSigFygShr2012 : (II.1) The division asymmetry $G(f)$ is directly fitted to data from tracking of single cell divisions. Then, this asymmetry $G(f)$ is incorporated into a partial differential equation model to optimize the proliferation rate $\gamma(A)$ based on cell size distributions measured in cell tracking experiments. The obtained rate $\gamma(A)$ is validated against independent data of cell division times in a stationary, postconfluent state. (II.2) The maximal migration velocity $v_{\text{max}}$, and consequently the migration rate $\mu$, is estimated from the time point of the transition to a stationary, postconfluent state and the average cell cycle time. (II.3) Finally, the growth rate $\alpha$ is optimized using simple Monte-Carlo-Simulations of independent cells based on the average cell cycle time using the already calibrated parameters $\gamma(A)$, $G(f)$. Note that the two mathematical models supporting the calibration are much simpler than the CA model, without spatial resolution or inter-cellular interaction.
II.1 Proliferation
@@
Figure 4: Distribution of cell sizes over time after transition to a stationary, postconfluent state (solid lines selected from experimental data (PulHufNevStrSigFygShr2012, , Fig. 4D)) is well described by partial differential equation (PDE Eq. (5), dashed lines), which only incorporates cell size-dependent proliferation $\gamma(A)$, Eq. (1), and asymmetric cell division $G(f)$, Eq. (4). Deviations from the experimental distributions originate mainly from overestimated tails at small cell sizes, which could hint to modification of the asymmetry of the cell division at cell sizes $<100\ \mu{}\text{m}^{2}$, see main text.
@@
$+\int\mathrm{d}f\ G(f)\ \gamma\left(\frac{A}{1-f}\right)\rho\left(\frac{A}{1-f},t\right)\ .$
The first term reflects the loss of cells with size $A$ due to division with proliferation rate $\gamma(A)$. The second and third term reflect the gain of cells with size $A$ due to division of larger cells with sizes $A/f$ or $A/(1-f)$, where one of the resulting daughter cells obtains a fraction $f$ or $1-f$, respectively, of the mother’s size. We solve Eq. (5) numerically by discretizing the cell size $A\to A_{i\in\mathbb{N}}$ and integrating the resulting set of ordinary equations with odeint Viretal2020 . We optimize the parameters $\gamma_{\text{max}}$, $A_{0}$, $m$ of the proliferation rate $\gamma(A)$ by minimizing the distance $\mathcal{D}$ between the distributions $\rho_{\text{exp}}(A_{i})$, $\rho_{\text{PDE}}(A_{i})$ obtained from experiment and PDE, respectively, at selected times $t$
@@
using the python package scipy.optimize Viretal2020 . We select experimental distributions at three time points from the experiment, where the distribution at the first time point $\rho_{\text{exp}}(A_{i},t\approx 0)$ serves as initial condition for the PDE while the other two time points $t\approx 0.5$ d and $t\approx 6$ d are the reference points for the optimization in Eq. (6).
Figure 5: Cell size-dependent proliferation rate $\gamma(A)$ (red line) optimized with PDE model based on cell area distributions after transition to a stationary, postconfluent state, see Fig. 4, is consistent with cell division times $\tau_{2}\approx\ln(2)/\gamma$ (black dots from single-cell tracking (PulHufNevStrSigFygShr2012, , Fig. 4E)). In particular, proliferation rates decrease rapidly for cell areas below $A<200\mu{}\text{m}^{2}$.
We obtain parameters $\gamma_{\text{max}}\approx 1.3$ d, $A_{0}\approx 80\text{ }\mu{}\text{m}^{2}$, $m\approx 3$, for which we observe good agreement both with the dynamics of the cell size distributions, see Fig. 4, and the division times of single cells, see Fig. 5. Deviations from the experimental distributions in Fig. 4 originate mainly from overestimated tails at small cell sizes, left of the maximum of each distribution. These deviations seem to be diminished, when the domain of the size fraction $p$, Eq. (4), is further narrowed around the symmetric case $p=1/2$. Thus, the deviations in Fig. 4 could hint at a narrowing of the asymmetry of the cell division at small at cell sizes $<100\ \mu{}\text{m}^{2}$, e.g., due to a minimal absolute size a cell must have. In fact, the distribution $G(p)$, which the PDE relies on, has been obtained for mother cell sizes $A\in[100,1000]\ \mu\text{m}^{2}$, due to the difficulty of tracking cells below $70\mu\text{m}^{2}$, while the cell size distributions $\rho(A)$ were mostly measured at $A<100\mu{}m^{2}$ PulHufNevStrSigFygShr2012 .
II.2 Migration
@@
with the average proliferation rate $\bar{\gamma}=\ln(2)/\tau_{2}$, given in terms of the reported average cell cycle time $\tau_{2}=0.75\pm 0.14$ d PulHufNevStrSigFygShr2012 . Note that in the context of the model, the rate $\bar{\gamma}$ should reflect the average proliferation rate over an ensemble of independently proliferating, Eq. (1), and growing, Eq. (3), cells, which is latter exploited for the calibration of (II.3) cell growth. Along with the initial culture size $A_{\text{cult}}(0)\approx 1.8\cdot 10^{4}\mu\text{m}^{2}$, the maximal migration velocity $v_{\text{max}}$ can then be estimated from the condition
@@
The range of the cell cycle time $\tau_{2}$ and the transition point $t_{T}=6\pm 1$ d imply a range for the maximal migration velocity $v_{\text{max}}\in[9,100]\ \mu\text{m}/\text{h}$. Note that the condition in Eq. (8) is a simplification and rather estimates a lower limit for $v_{\text{max}}$. The transition from exponential to subexponential growth is expected before the rim of the cell culture actually reaches an outward velocity $v_{\text{max}}$. For instance, at $t=5$ d the culture already exhibits deviations from exponential growth, while according to Eq. (8) the velocity of the boundary at this point is only $8-30\ \mu$m. In order to reflect the estimated range for $v_{\text{max}}$, we use its upper limit $v_{\text{max}}=100\ \mu\text{m}/\text{h}$ in the following, but also report the results for $v_{\text{max}}=50\ \mu\text{m}/\text{h}$ in SI Figs. 12-15.
II.3 Cell growth
The average proliferation rate $\bar{\gamma}$ derived from the measured average cell cycle time is also used to estimate the cell growth rate $\alpha$: Firstly, the maximal cell size is set to $A_{\text{max}}=990\ \mu$m based on tracking of sizes of single cells (PulHufNevStrSigFygShr2012, , Fig. 4A) and the observed minimal cell density. For the phase of exponential growth of the culture, it is reasonable to assume that cells proliferate and grow independently of each other according to Eqs. (1), (3), (4). We compute these simple dynamics numerically for ensembles of cells using Monte-Carlo simulations. In particular, we use the parameters of proliferation $\gamma(A)$, $G(p)$ as calibrated above and additionally optimize the growth rate $\alpha$ of logistic cell growth Eq. (3) such that the resulting exponential slope of the total size of the cell ensemble matches with the slope $\bar{\gamma}=\ln(2)/\tau_{2}$ corresponding to the average cell cycle time $\tau_{2}=0.75\pm 0.14$ d measured independently by single-cell tracking. From this optimization a growth rate $\alpha=1.3\pm 0.3\ \text{d}^{-1}$ is obtained.
[…]
III Results
@@
Figure 7: Emergent dynamics of calibrated CA model (red line) reproduces experimentally observed cell densities of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 1C), corresponding data to areas displayed in Fig. 6). The cell density is initially constant (low-density, confluent regime) and increases rapidly around time $t_{T}=6\pm 1$ d (transition to a stationary, postconfluent state) implying dense packing of cells, which become smaller and smaller due to slowing but continued proliferation without cell growth. Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region. Model parameters calibrated as written in the main text.
We implement the previously calibrated parameters into the CA model and compare the emerging dynamics of culture size and cell density with the experiment of epithelial colony growth, see Figs. 6 and 7. In the CA, the culture size is computed as the total area of all grid points containing at least a single cell and the cell density is computed as ratio between this area and the total number of cells. The resulting dynamics of culture size and cell density match with the experimental ones. In particular, the CA model does not only quantitatively reproduce the exponential growth and constant density in the initial low-density confluent regime and the time point of the transition to a stationary, postconfluent state, but also the curve at and after this transition. The predictions of the CA model are also consistent for the case of initially uniformly seeded cells, see Fig. 8, which reflects the regime directly at and long after the transition to a stationary, postconfluent state. Furthermore, the transition is accompanied by a rapid decrease in the emergent cell velocities, which are comparable between CA model and experiment, see Fig. 9. In particular, there is a velocity gradient from the inside to the outside of the colony and an outwardly biased flow, see Fig. 10.
@@
The match of the emergent features, i.e., dynamics of culture size and cell density, between experiment and CA model suggests that they are determined by the single-cell dynamics. In particular, the CA model does not incorporate any inter-cellular interactions, besides the exclusion principle. In the CA model, the transition of the dynamics at $t_{T}$ originates solely from the limitation that the boundary of the culture cannot expand faster outwards than the maximal migration velocity $v_{\text{max}}$ of the cells. Beyond the corresponding critical size of the culture, cells in the center of the culture no longer get access to the necessary space to grow, i.e. $\mathrm{d}A/\mathrm{d}t\to 0$, and consequently become smaller due to continued proliferation. The latter gets slower due to the size-dependent, sigmoidal proliferation rate $\gamma(A)$. The capacity of cells to grow separates their proliferation behavior into a pre- ($A\gg A_{0}\to\gamma(A)=\gamma_{\text{max}}$) and posttransition regime ($A\sim A_{0}\to\gamma(A)\ll\gamma_{\text{max}}$) according to Eq. (1). Note that based on the assumption that for times $t\gg t_{T}$ effectively only cells in a boundary layer with thickness $\Delta R$ of the culture can proliferate with rate $\bar{\gamma}$, the asymptotic, posttransition growth may be estimated by $\mathrm{d}A/\mathrm{d}t\approx 2\sqrt{\pi}\bar{\gamma}\Delta R\cdot\sqrt{A}=\kappa\sqrt{A}$ or $A(t\gg t_{T})=\kappa^{2}/4\cdot(t-t_{0})^{2}$, analogous to recent estimates for spherical cultures FraMicAlaKunVosLan2023 , while $\Delta R(v_{\text{max}})$ may depend on the maximal migration velocity. We obtain an effective thickness of the proliferating rim $\Delta R=370\pm 10\ \mu$m ($t_{0}=3.3\pm 0.2$ d) from the CA simulation, consistent with rescaled results from a previous theoretical model of epithelial colonies AlaHatLowVoi2015 .
@@
Figure 8: Emergent dynamics of calibrated CA model (red line) consistent with experimentally observed cell size median for initially uniformly seeded cells (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 3C), transition $t_{T}$ to a stationary, postconfluent state corresponds roughly to $1.5$ d). Cell size median initially drops rapidly and then saturates at small cell sizes $A\sim 40\ \mu\text{m}^{2}$. Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region.
@@
Figure 9: Emergent dynamics of calibrated CA model (red line) reproduces qualitatively the experimentally observed cell velocities of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 2C), corresponding data to Figs. 6 and 7). The average cell velocity is initially constant (low-density, confluent regime) and decreases rapidly around time $t_{T}=6\pm 1$ d (transition to a stationary, postconfluent state). Presented are the root-mean-square velocities, where the model velocities are computed from a single simulation run over time increments of $580$ min to neglect small-scale oscillations. Due to the presence of a radial velocity gradient the corresponding standard deviation of the velocities (red shaded region) is quite large.
@@
Figure 10: CA model displays an outwards biased cell motion (right panel) accompanied by a velocity gradient from inner region of colony to its periphery (left and right panel encoded by color). Displayed are cell velocities $\bar{v}$ of the last $580$ min of the simulation (posttransition) spatially averaged (left panel: $4\times 4$ grid squares, right panel: higher resolution $2\times 2$ grid squares) where arrows represent direction of motion (left panel) and color magnitude of velocity (linear scale in left panel and logarithmic scale in right panel for visibility, missing arrows or white squares mean no measurable cell motion).
@@
In the CA model, the observed rapid decrease in cell velocity at the transition results from cells in the inner region of the colony slowing down due to insufficient space to migrate, which generates a velocity gradient from inside to outside, see Fig. 10. At the same time cell motion transitions from undirected to a collective outwards bias, see also Fig. 11. This outwards directed motion is an emergent feature of the interplay between the migration rate, the exclusion principle, and the fact that more free space is available outwards. It should be emphasized that for the CA model the maximal migration velocity $v_{\text{max}}$ is significantly bigger than the velocities that are measured for individual cells inside the culture, see Fig. 9, due to the exclusion principle. For instance, 1 to 3 occupied neighboring grid squares reduce the actual migration velocity by a factor of $3/4$ to $1/4$. Moreover, the migration velocity in the CA model represents the speed of instantaneous, undirected movement of single cells, whereas in in-vitro experiments typically collective, directed movement of groups of cells is measured, e.g., directed outwards in case of an expanding culture. Consequently, this directed motion exhibits much smaller velocities, consistent with previous experimental measurements of cell velocity in epithelial tissue $\sim 25\mu\text{m}/\text{h}$ depending on the setting (wound-cut technique, expanding colony, …) PulHufNevStrSigFygShr2012 ; TauChuSaiSat1979 ; RosMis1980 ; PetRefGraPouLadBugSil2010 ; JenPedMorParPedNej2019 . Note that the exact value of the average velocity depends to some degree on the chosen definition, i.e., the chosen time increment between two cell position measurements and whether the average is taken over the whole colony or a specific location. Nevertheless, we can infer from the match of simulation and experiment in Fig. 9 that the course of the cell velocities, initially constant with a rapid decrease around the transition point, is reproduced by the CA model.
IV Discussion

### v2

@@ III Parameter calibration
Figure 3: Asymmetry $G(f)$ of the daughter cell sizes after cell division is well described by a Gaussian distribution with expectation value $\bar{f}=1/2$, standard deviation $\sigma=1/8$, and a cut-off at large and small cell sizes $f\in[1/4,3/4]$, reflecting the absence of smaller or larger daughter cells (black dots represent data from Ref. (PulHufNevStrSigFygShr2012, , Fig. 4C (inset)) obtained from tracking single-cell divisions, Gaussian distribution shown as red line). It is assumed that during division of a cell, one daughter cell has a fraction $f$ of the size of the mother cell, while the other cell has a fraction $(1-f)$.
@@
The biologically motivated parameters $\gamma(A)$, $v_{\text{max}}$, $\alpha$ of the CA model are calibrated based on in-vitro observations, mostly of single-cell dynamics, during the growth of normally differentiating epithelial cell cultures of Madin-Darby canine kidney cells PulHufNevStrSigFygShr2012 : (III.1) The division asymmetry $G(f)$ is directly fitted to data from tracking of single cell divisions. Then, this asymmetry $G(f)$ is incorporated into a partial differential equation model to optimize the proliferation rate $\gamma(A)$ based on cell size distributions measured in cell tracking experiments. The obtained rate $\gamma(A)$ is validated against independent data of cell division times in a stationary, postconfluent state. (III.2) The maximal migration velocity $v_{\text{max}}$, and consequently the migration rate $\mu$, is estimated from the time point of the transition to a stationary, postconfluent state and the average cell cycle time. (III.3) Finally, the growth rate $\alpha$ is optimized using simple Monte-Carlo-Simulations of independent cells based on the average cell cycle time using the already calibrated parameters $\gamma(A)$, $G(f)$. Note that the two mathematical models supporting the calibration are much simpler than the CA model, without spatial resolution or inter-cellular interaction.
III.1 Proliferation
@@
Figure 4: Distribution of cell sizes over time after transition to a stationary, postconfluent state (solid lines selected from experimental data (PulHufNevStrSigFygShr2012, , Fig. 4D)) is well described by partial differential equation (PDE Eq. (5), dashed lines), which only incorporates cell size-dependent proliferation $\gamma(A)$, Eq. (1), and asymmetric cell division $G(f)$, Eq. (4). The distribution at day 0 serves as initial condition for the PDE. Deviations from the experimental distributions originate mainly from overestimated tails at small cell sizes, which could hint to modification of the asymmetry of the cell division at cell sizes $<100\ \mu{}\text{m}^{2}$, see main text.
@@
$+\int\mathrm{d}f\ G(f)\ \gamma\left(\frac{A}{1-f}\right)\rho\left(\frac{A}{1-f},t\right)$
=$\underset{G(f)=G(1-f)}{=}$ $-\gamma(A)\rho(A,t)+2\int\mathrm{d}f\ G(f)\ \gamma\left(\frac{A}{f}\right)\rho\left(\frac{A}{f},t\right)\ .$
The first term reflects the loss of cells with size $A$ due to division with proliferation rate $\gamma(A)$. The second and third term reflect the gain of cells with size $A$ due to division of larger cells with sizes $A/f$ or $A/(1-f)$, where one of the resulting daughter cells obtains a fraction $f$ or $1-f$, respectively, of the mother’s size. Note that Eq. (5) simplifies to the equation proposed in Ref. PulHufNevStrSigFygShr2012 , if a delta distribution $G(f)=\delta(f-1/2)$ is assumed. We solve Eq. (5) numerically by discretizing the cell size $A\to A_{i\in\mathbb{N}}$ and integrating the resulting set of ordinary equations with odeint Viretal2020 . We optimize the parameters $\gamma_{\text{max}}$, $A_{0}$, $m$ of the proliferation rate $\gamma(A)$ by minimizing the distance $\mathcal{D}$ between the distributions $\rho_{\text{exp}}(A_{i})$, $\rho_{\text{PDE}}(A_{i})$ obtained from experiment and PDE, respectively, at selected times $t$
@@
using the python package scipy.optimize Viretal2020 . We select experimental distributions at three characteristic time points from the experiment, where the distribution at the first time point $\rho_{\text{exp}}(A_{i},t\approx 0)$ serves as initial condition for the PDE while the other two time points $t\approx 0.5$ d and $t\approx 6$ d are the reference points for the optimization in Eq. (6). While distributions are reported at more intermediate time points in Ref. PulHufNevStrSigFygShr2012 , they do not significantly contribute to the optimization as these distributions transition smoothly into each other and thus can be neglected for clarity.
Figure 5: Cell size-dependent proliferation rate $\gamma(A)$ (red line, Eq. (1) with $\gamma_{\text{max}}=1.3\ \text{d}^{-1}$, $A_{0}=80\text{ }\mu{}\text{m}^{2}$, $m=3$) optimized with PDE model based on cell area distributions after transition to a stationary, postconfluent state, see Fig. 4, is consistent with cell division times $\tau_{2}\approx\ln(2)/\gamma$ (black dots from single-cell tracking (PulHufNevStrSigFygShr2012, , Fig. 4E)). In particular, proliferation rates decrease rapidly for cell areas below $A<200\mu{}\text{m}^{2}$.
We obtain parameters $\gamma_{\text{max}}\approx 1.3\ \text{d}^{-1}$, $A_{0}\approx 80\text{ }\mu{}\text{m}^{2}$, $m\approx 3$, for which we observe good agreement both with the dynamics of the cell size distributions, see Fig. 4, and the division times of single cells, see Fig. 5. Deviations from the experimental distributions in Fig. 4 originate mainly from overestimated tails at small cell sizes, left of the maximum of each distribution. These deviations seem to be diminished, when the domain of the size fraction $p$, Eq. (4), is further narrowed around the symmetric case $p=1/2$. Thus, the deviations in Fig. 4 could hint at a narrowing of the asymmetry of the cell division at small at cell sizes $<100\ \mu{}\text{m}^{2}$, e.g., due to a minimal absolute size a cell must have. In fact, the distribution $G(p)$, which the PDE relies on, has been obtained for mother cell sizes $A\in[100,1000]\ \mu\text{m}^{2}$, due to the difficulty of tracking cells below $70\ \mu\text{m}^{2}$, while the cell size distributions $\rho(A)$ were mostly measured at $A<100\mu{}m^{2}$ PulHufNevStrSigFygShr2012 . Furthermore, the PDE assumes conservation of total cell area for simplicity, which is not always the case in the experiment (PulHufNevStrSigFygShr2012, , Fig. 4c, lower inset).
III.2 Migration
@@
with the average proliferation rate $\bar{\gamma}=\ln(2)/\bar{\tau}_{2}$, given in terms of the reported average cell cycle time $\bar{\tau}_{2}=0.75\pm 0.14$ d PulHufNevStrSigFygShr2012 . Note that in the context of the model, the rate $\bar{\gamma}$ should reflect the average proliferation rate over an ensemble of independently proliferating, Eq. (1), and growing, Eq. (3), cells, which is latter exploited for the calibration of (III.3) cell growth. Along with the initial culture size $A_{\text{cult}}(0)\approx 1.8\cdot 10^{4}\mu\text{m}^{2}$, the maximal migration velocity $v_{\text{max}}$ can then be estimated from the condition
@@
The range of the cell cycle time $\bar{\tau}_{2}$ and the transition point $t_{T}=6\pm 1$ d imply a range for the maximal migration velocity $v_{\text{max}}\in[9,100]\ \mu\text{m}/\text{h}$. In Ref. PulHufNevStrSigFygShr2012 a coarse consistency check between the observed average cell velocity, transition time $t_{T}$, average cell cycle time $\bar{\tau}_{2}$, and colony area $A_{\text{cult}}(t_{T})$ was performed based on the same assumption as Eq. (8). However, the ranges for $t_{T}$ and $\bar{\tau}_{2}$ were neglected and the maximal and average cell velocity were assumed to be the same, while the former is a single-cell parameter and the latter an emergent feature from the interaction with other cells. Note that the condition in Eq. (8) is a simplification and rather estimates a lower limit for $v_{\text{max}}$. The transition from exponential to subexponential growth is expected before the rim of the cell culture actually reaches an outward velocity $v_{\text{max}}$. For instance, at $t=5$ d the culture already exhibits deviations from exponential growth, while according to Eq. (8) the velocity of the boundary at this point is only $8-30\ \mu$m. In order to reflect the estimated range for $v_{\text{max}}$, we use its upper limit $v_{\text{max}}=100\ \mu\text{m}/\text{h}$ in the following, but also report the results for $v_{\text{max}}=50\ \mu\text{m}/\text{h}$ in SI Figs. 4-7.
III.3 Cell growth
The average proliferation rate $\bar{\gamma}$ derived from the measured average cell cycle time is also used to estimate the cell growth rate $\alpha$: Firstly, the maximal cell size is set to $A_{\text{max}}=990\ \mu\text{m}^{2}$ based on tracking of sizes of single cells (PulHufNevStrSigFygShr2012, , Fig. 4A) and the observed minimal cell density. For the phase of exponential growth of the culture, it is reasonable to assume that cells proliferate and grow independently of each other according to Eqs. (1), (3), (4). We compute these simple dynamics numerically for ensembles of cells using Monte-Carlo simulations. In particular, we use the parameters of proliferation $\gamma(A)$, $G(p)$ as calibrated above and additionally optimize the growth rate $\alpha$ of logistic cell growth Eq. (3) such that the resulting exponential slope of the total size of the cell ensemble matches with the slope $\bar{\gamma}=\ln(2)/\bar{\tau}_{2}$ corresponding to the average cell cycle time $\bar{\tau}_{2}=0.75\pm 0.14$ d measured independently by single-cell tracking. From this optimization a growth rate $\alpha=1.3\pm 0.3\ \text{d}^{-1}$ is obtained.
III.4 Initial and boundary conditions
For expanding culture experiments, the size of the spatial grid is chosen sufficiently large, such that cells do not reach the boundary of the grid during simulation. We obtain the initial condition of the CA model by starting from a single cell of maximal size in the center of the grid and then simulating the dynamics until the size of the culture matches the initial one in the experiment. For experiments where cells are seeded homogeneously, a $12\times 12$-grid with reflecting boundary conditions is used and as initial condition a single cell of maximal size is placed in every second grid point, reflecting the experimental seeding density of $600$ cells/mm2. Finally, the time increment in the CA is set sufficiently small $\Delta t=6$ min such that $\Delta t\ll\text{max}\{\gamma_{\text{max}}^{-1},\mu^{-1}\}$ and $\mathrm{\Delta}{}A/A\ll 1$.
IV Results
@@
Figure 7: Emergent dynamics of calibrated CA model (red line) reproduces experimentally observed cell densities of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 1C), corresponding data to areas displayed in Fig. 6). The cell density is initially constant (low-density, confluent regime) and increases rapidly around time $t_{T}=6\pm 1$ d (transition to a stationary, postconfluent state) implying dense packing of cells, which become smaller and smaller due to slowing but continued proliferation without cell growth. Note that the small variations (±$\pm 30\%$) until day 6 appear larger due to the logarithmic y-scale and were interpreted as constant density by the original experimenter PulHufNevStrSigFygShr2012 . Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region. Model parameters calibrated as written in the main text.
We find quantitative agreement between the calibrated CA model and independent experiments with respect to several emerging features of the epithelial cell culture. This validates our minimal model and showcases its power to predict emergent tissue wide dynamics from single cell behavior. The agreement also suggests that the exclusion principle as sole inter-cellular interaction paired with a size-dependent proliferation rate of each cell is sufficient to explain the emerging dynamics of contact inhibition:
We implement the previously calibrated parameters into the CA model and compare the emerging dynamics of culture size and cell density with the experiment of epithelial colony growth, see Figs. 6 and 7. In the CA, the culture size is computed as the total area of all grid points containing at least a single cell and the cell density is computed as ratio between this area and the total number of cells. The resulting dynamics of culture size and cell density match with the experimental ones. In particular, the CA model does not only quantitatively reproduce the exponential growth and constant density in the initial low-density confluent regime and the time point of the transition to a stationary, postconfluent state, but also the curve at and after this transition. The predictions of the CA model are also consistent for the case of initially homogeneously seeded cells, see Fig. 8, which reflects the regime directly at and long after the transition to a stationary, postconfluent state. Furthermore, the transition is accompanied by a rapid decrease in the emergent cell velocities, which are comparable between CA model and experiment, see Fig. 9. In particular, there is a velocity gradient from the inside to the outside of the colony and an outwardly biased flow, see Fig. 10 as well as SI Figs. 1 and 2.
@@
The match of the emergent features, i.e., dynamics of culture size and cell density, between experiment and CA model suggests that they are determined by the single-cell dynamics. In particular, the CA model does not incorporate any inter-cellular interactions, besides the exclusion principle. In the CA model, the transition of the dynamics at $t_{T}$ originates solely from the limitation that the boundary of the culture cannot expand faster outwards than the maximal migration velocity $v_{\text{max}}$ of the cells. Beyond the corresponding critical size of the culture, cells in the center of the culture no longer get access to the necessary space to grow, i.e. $\mathrm{d}A/\mathrm{d}t\to 0$, and consequently become smaller due to continued proliferation. The latter gets slower due to the size-dependent, sigmoidal proliferation rate $\gamma(A)$. The capacity of cells to grow separates their proliferation behavior into a pre- ($A\gg A_{0}\to\gamma(A)=\gamma_{\text{max}}$) and posttransition regime ($A\sim A_{0}\to\gamma(A)\ll\gamma_{\text{max}}$) according to Eq. (1). Note that based on the assumption that for times $t\gg t_{T}$ effectively only cells in a boundary layer with thickness $\Delta R$ of the culture can proliferate with rate $\bar{\gamma}$, the asymptotic, posttransition growth may be estimated by $\mathrm{d}A/\mathrm{d}t\approx 2\sqrt{\pi}\bar{\gamma}\Delta R\cdot\sqrt{A}=\kappa\sqrt{A}$ or $A(t\gg t_{T})=\kappa^{2}/4\cdot(t-t_{0})^{2}$, analogous to recent estimates for spherical cultures FraMicAlaKunVosLan2023 , while $\Delta R(v_{\text{max}})$ may depend on the maximal migration velocity. We obtain an effective thickness of the proliferating rim $\Delta R=370\pm 10\ \mu$m ($t_{0}=3.3\pm 0.2$ d) from the CA simulation, consistent with rescaled results from a previous theoretical model of epithelial colonies AlaHatLowVoi2015 . Note that while the cell size-dependent division times $t_{2}$ exhibited by the calibrated CA model closely resemble the experimental ones, see SI Fig. 3, the distribution of these times is broader than in the experiment, see inset in Fig. 3. In fact, the division times in the CA model follow an exponential distribution by design for mathematical and numerical convenience, whereas biological cells divide at more regular intervals leading typically to a narrower Erlang-like distribution. However, our results imply that this difference in the cell cycle distributions does not affect the emergent dynamics of the colony and the implementation of cell division according to regular cell cycles into the CA model is straightforward.
@@
Figure 8: Emergent dynamics of calibrated CA model (red line) consistent with experimentally observed cell size median for initially homogeneously seeded cells (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 3C), transition $t_{T}$ to a stationary, postconfluent state corresponds roughly to $1.5$ d). Cell size median initially drops rapidly and then saturates at small cell sizes $A\sim 40\ \mu\text{m}^{2}$. Results from CA model are averaged over $15$ independent runs and standard deviation is displayed as red shaded region.
@@
Figure 9: Emergent dynamics of calibrated CA model (red line) reproduces qualitatively the experimentally observed cell velocities of expanding epithelial colony (black dots from Ref. (PulHufNevStrSigFygShr2012, , Fig. 2C), corresponding data to Figs. 6 and 7). The average cell velocity is initially constant (low-density, confluent regime) and decreases rapidly around time $t_{T}=6\pm 1$ d (transition to a stationary, postconfluent state). Presented are the root-mean-square velocities, where the model velocities are computed from a single simulation run over time increments of $580$ min to neglect small-scale oscillations. Due to the presence of a radial velocity gradient, see Fig. 10, the corresponding standard deviation of the velocities (red shaded region) is quite large.
@@
Figure 10: CA model displays an outwards biased cell motion (right panel) accompanied by a velocity gradient from inner region of colony to its periphery (left and right panel encoded by color). Displayed are cell velocities $\bar{v}$ of the last $580$ min of the simulation (posttransition) spatially averaged (left panel: $4\times 4$ grid squares, right panel: higher resolution $2\times 2$ grid squares) where arrows represent direction of motion (left panel) and color magnitude of velocity (linear scale in left panel and logarithmic scale in right panel for visibility, missing arrows or white squares mean no measurable cell motion). For a radial profile of the velocities see SI Fig. 2.
@@
In the CA model, the observed rapid decrease in cell velocity at the transition results from cells in the inner region of the colony slowing down due to insufficient space to migrate, which generates a velocity gradient from inside to outside, see Fig. 10. At the same time cell motion transitions from undirected to a collective outwards bias, see also Fig. 1. This outwards directed motion is an emergent feature of the interplay between the migration rate, the exclusion principle, and the fact that more free space is available outwards. It should be emphasized that for the CA model the maximal migration velocity $v_{\text{max}}$ is significantly bigger than the velocities that are measured for individual cells inside the culture, see Fig. 9, due to the exclusion principle. For instance, 1 to 3 occupied neighboring grid squares reduce the actual migration velocity by a factor of $3/4$ to $1/4$. Moreover, the migration velocity in the CA model represents the speed of instantaneous, undirected movement of single cells, whereas in in-vitro experiments typically collective, directed movement of groups of cells is measured, e.g., directed outwards in case of an expanding culture. Consequently, this directed motion exhibits much smaller velocities, consistent with previous experimental measurements of cell velocity in epithelial tissue $\sim 25\mu\text{m}/\text{h}$ depending on the setting (wound-cut technique, expanding colony, …) PulHufNevStrSigFygShr2012 ; TauChuSaiSat1979 ; RosMis1980 ; PetRefGraPouLadBugSil2010 ; JenPedMorParPedNej2019 . Note that the exact value of the average velocity depends to some degree on the chosen definition, i.e., the chosen time increment between two cell position measurements and whether the average is taken over the whole colony or a specific location. In Ref. PulHufNevStrSigFygShr2012 the cell velocity was extracted from four local $450\times 336\ \mu\text{m}^{2}$ fields of view, while for the CA model it is possible to average over the entire cell culture. This explains the large range of velocities observed in the CA model, which reflects the radial velocity gradient, see Fig. 10, while the standard error for the experimental velocities is not reported. Nevertheless, we can infer from the match of simulation and experiment in Fig. 9 that the course of the cell velocities, initially constant with a rapid decrease around the transition point, is reproduced by the CA model.
V Discussion